OpenStudy (anonymous):

You kick a soccer ball straight up into the air with a speed of 21.2 m/s. How long does it take the soccer ball to reach its highest point?

OpenStudy (anonymous):

\[v=v(0) + at\] at the highest point, the velocity will be 0, the initial velocity is 21.2, and a is 9.81. plug in to get t=21.2/9.81 = 2.16 seconds

OpenStudy (anonymous):

You're conducting a physics experiment on another planet. You drop a rock from a height of 2.3 m and it hits the ground 1.1 seconds later. What is the acceleration due to gravity on this planet?

OpenStudy (theeric):

If your professor is picky about showing your work, you'll want to change that just a little bit to use the correct signs (+-).

OpenStudy (theeric):

That is a different question... Why don't you use a different post, @bigsatto ?

OpenStudy (anonymous):

quicker responses @theEric

OpenStudy (anonymous):

h = (1/2) a t^2 2.3 = (1/2) a (1.1)^2 you are nearly done.